OpenStudy (anonymous):

help i am stuck on this homework problem i know how to do the formula i just get stuck on how to finish it.... find

OpenStudy (anonymous):



OpenStudy (anonymous):

whenever two lines like this intersect: < abc = < dbe and < abd = < cbe this means that 6x - 7 = 4x + 23 solve for x then plug it in.

OpenStudy (anonymous):

i know that but this is what i get when i do that 

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

that's fine. this means < abc = 83.

OpenStudy (anonymous):

how do u get 83

OpenStudy (anonymous):

@Euler271

Directrix (directrix):

The measure of < ABC is given algebraically as (6x - 7)/ Assuming x = 15, then the measure of <ABC = 6*15 - 7 = ? @leroysgurl15

OpenStudy (anonymous):

gotcha thanks

Directrix (directrix):

You are welcome. Also, note the reason that those two angles are congruent is that they are vertical angles. By a theorem, vertical angles are congruent and therefore have the same measure.